Find the coordinates for the midpoint of the segment with endpoints given.
(5, 6) and (8, 2)


Answer by timvanswearingen(106)

Your midpoint is another ordered pair (x,y)
Add up the x-coordinates and divide by 2, add up the y-coordinates and divide by 2.
x:(5+8)/2 = 13/2
y:(6+2)/2 = 8/2 = 4
So the midpoint is (13/2 , 4)